# Average-Case Local Computation Algorithms

### v1

@@ 1.2 Our Results: Joint Sampling
A Local Access Generator (LAG) of a random object $G$ sampled from a distribution $\mathcal{G}$, is an oracle that provides access to $G$ by answering various types of supported queries, given a sequence of random bits $\vec{r}$. We say the LAG is memoryless if it does not store its answers to prior queries. We require that the LAG satisfies the following:
  • Consistency. The responses of the local-access generator to all probes throughout the entire execution must be consistent with a single object $G$.
  • Distribution equivalence. The distribution $\mathcal{G}^{\prime}$ sampled by the LAG must be within $n^{-c}$ from $\mathcal{G}$ in TV distance, for any desired constant $c$.
Again, as in the case of LCAs, we desire Local Access Generators to be as efficient as possible per query. We also strongly desire the LAG to be memoryless (a requirement in the setting of LCAs, but not always achieved for LAGs), and our result achieves this goal.
@@
We show that such an approach is also fruitful in the setting of LCAs. We again focus on the dense Erdős-Rényi model, and this time on the extensively studied problem [Gha16, LRY17, GU19, Gha22] of Maximal Independent Set (MIS).
The frontier result of Ghaffari [Gha22] provides an LCA for MIS with per-query runtime $\operatorname{poly}(\Delta\log n)$, and a local sampling implementation of dense Erdős-Rényi graphs is straightforward. However, composing these algorithms does not give a sublinear runtime. Our result achieves runtime $\operatorname{polylog}(n)$ for $p\geq 1/\operatorname{polylog}(n)$ per query, both for queries to the random graph and to its MIS:
@@
Next, we overview the proofs of these results.
1.3 Proof Overview: Spanners for Erdős-Rényi
@@
We recall the statement of the result. See 1.6 For polynomially dense Erdős-Rényi graphs, our connectivity rule is as follows. We designate a sublinear-size set of vertices in $G$ as centers, which we denote $\mathcal{C}$. We then retain in $H$ all edges between centers. Finally, every non-center vertex adds the first edge from itself to $\mathcal{C}$ in $H$.
By choosing the size of $\mathcal{C}$ appropriately, we ensure that the following three conditions hold with high probability:
  • there are $o(n)$ intra-center edges,
  • every non-center vertex has an edge to the center with high probability,
  • the center has constant diameter.
The first condition holds as we choose the size of the center sufficiently small such that $|\mathcal{C}|^{2}p=o(n)$. The second holds as we choose the size of center sufficient large such that $|\mathcal{C}|p=\Omega(\log n)$, so every vertex is incident to at least one center whp. To see why the first con final condition holds, observe that $\mathcal{C}$ is itself distributed as $G(|\mathcal{C}|,p)$, so we can appeal to known results on the diameter of $G(m,p)$.
To implement this connectivity rule as an LCA, we break into the sparse case (where the vertex simply queries its entire adjacency list and chooses the least-ranked edge to keep) or the dense case (where the vertex queries the adjacency matrix until it finds the first edge). If we additionally assume that the adjacency list of each vertex is sorted in ascending order, we can perform this check in constant time.
To modify this algorithm to have constant query time (Theorem 1.7), even when we do not have a sorted adjacency list, we change the connectivity rule such that each vertex randomly keeps $O(\log n)$ edges from itself to the center in expectation. This does not affect the center, which retains constant stretch, and allows us to decide connectivity only examining random bits relevant to the particular queried edge. This does increase the edge count in $H$ to $O(n\log n)$, but we maintain constant stretch. For the formal proofs, see Section 3.
To obtain even smaller stretch for sufficiently dense graphs, we modify the connectivity rule. See 1.8 We again let the first $T$ vertices be centers, for appropriately chosen $T$. In this case however, our graph is sufficiently dense that with high probability every pair of centers is connected through a non-center vertex, providing a path of length $2$ between them in $G$. We then retain the lexicographically first path between every pair of centers, and an edge from every non-center vertex to its first center. By doing so we obtain a stretch of $4$. Building a fast local implementation of this algorithm involves first subsampling the graph to have degree barely high enough to preserve connectivity using the rule, then arguing we can locally search in this pre-sparsified graph. For the formal proof, see Section 3.
1.4 Proof Overview: Sparse Spanning Subgraphs for Erdős-Rényi
Here we are given probe access to $G\leftarrow G(n,p)$ and wish to provide an LCA which gives a user access to a sparse connected subgraph $H\subseteq G$ with very few edges. See 1.9 Without essential loss of generality, we assume the edge probability is exactly $p^{*}=7\log n/n$ (as otherwise we can use the idea of Observation 1.4 to subsample as a first step).
@@ 1.4 Proof Overview: Sparse Spanning Subgraphs for Erdős-Rényi
We overview the needed properties below.
  • Connectivity. The connectivity argument is the most delicate of the paper. We define a set of events $\mathcal{E}$ that partition the space of possible graphs, and denote a subset of events $\mathcal{E}_{G}\subset\mathcal{E}$ as good. We first show that a random graph lies in a good event with high probability. Next, we show that for every good event $E\in\mathcal{E}_{G}$, sampling a random $G$ that satisfies $E$ results in a graph that the algorithm succeeds on (in fact, we prove this with high probability over $G$).
    Each such event specifies the presence of absence of a subset of edges in the graph. At a high level, these specifications capture the view of the algorithm up to the point that the leader vertices select their administrators. We define good events as those in which all administrators have many bits of entropy remaining in their neighborhoods, which allows us to argue they maintain connectivity with high probability.
@@
  • Local Implementation. One can see that the algorithm constitutes a $4$-round distributed algorithm, and hence can be implemented in per-query work $O(\Delta^{4})$ via the reduction of Parnas and Ron [PR07]. However, we note that we can first subsample the graph $G$ to have edge probability $p^{*}=7\log n$ (which we do in a global fashion using the random tape of the LCA). By Observation 1.4, this produces whp a connected subgraph of $G$ that is itself distributed $G(n,p^{*})$. Subsequently, in each distributed round we explore only the neighbors of $v$ that are retained in the subsampled graph, resulting in total work $O(\Delta\log^{4}n)$.
1.5 Proof Overview: Spanners for Preferential Attachment
@@
We now build low-stretch spanning trees for preferential attachment graphs with sufficiently high degree parameter $\mu\geq c_{\mu}\log n$. See 1.10 For clarity, in this section we let $v_{i}$ for $i\in[n]$ denote the vertex that was added in round $i$. For the formal proof, see Section 5. Our algorithm is a straightforward consequence of a structural result about preferential attachment graphs. We show that with high probability, every vertex $v$ either has an edge to the highest-degree vertex, or has an edge to a vertex with substantially higher degree:
Lemma 1.13.
To show Lemma 1.13, we prove that there is a global constant $c$ such that with high probability:
  1. 1.
    every $v_{i}$ with $i\geq c$ has a neighbor with degree at least $2d_{v_{i}}$,
  2. 2.
    every $v_{i},v_{j}$ for $i,j\leq c$ are connected.
The second item is a simple consequence of our choice of $\mu$. The first we show as follows. First, whp every vertex has an edge to a vertex to significantly lower index, and second, whp vertices of significantly lower index have significantly higher degree. For both cases, we rely on a tail bound for the degrees of Preferential Attachment graphs established by [DKR18].
@@
Given this lemma, Theorem 1.10 follows via a simple connectivity rule: every vertex retains an edge to its highest-degree neighbor. Then a simple potential argument (that the degree cannot increase by a factor of $2$ more than $\log(m)$ times) establishes that we obtain a spanner $O(\log(d_{\max}))=O(\log n)$ stretch. Furthermore, we show that the worst-case runtime is $O(\sqrt{n})$ and the average-case runtime over all possible queries is $\operatorname{polylog}(n)$. To see the size bound, note that every vertex that is not of globally highest degree adds exactly one edge to $H$.
1.6 Proof Overview: Joint Sampling
@@
Finally, we describe our algorithm that provides access to a random graph together with its MIS. See 1.12 We describe the sampling algorithm in a global fashion, and then describe how to implement it via an LCA. We gradually grow the MIS $M$, by instantiating $M=(1)$ where $1$ is the first vertex, and sequentially determining the smallest vertex that is not connected to $M$, and add it to $M$, continuing until we exhaust the vertex set. For a fixed $M=(v_{1},\ldots,v_{t})$, each subsequent vertex $u>v_{t}$ is not connected to any element of $M$ in $G(n,p)$ with probability $(1-p)^{|M|}$. For a fixed $u$, we can determine if $(v_{i},u)$ is in $G$ for every $i$ by simply sampling the edge, and thus determine if $u$ should be added to $M$. However, this procedure would take linear time to determine the MIS. Instead, we use local sampling of the Geometric distribution to find the next element of $M$. Once we sample $r\sim\mathrm{Geom}(1-p)$, we let the next element of $M$ be $u=v_{t}+r$. For all $u^{\prime}\in(v,u)$, this virtually conditions on the event that at least one of $(u,v_{i})$ is present in the graph for $i\leq t$.
@@ 2 Preliminaries
We formally define the Erdős-Rényi and Preferential Attachment graph models. Note that in all proofs, $\Gamma(v)=\Gamma_{G}(v)$ refers to the neighbors of $v$ in the original graph $G$ that the LCA has probe access to. We will always denote the original graph as $G$, and subgraphs or spanners as $H$.
@@
Definition 2.2 (Preferential Attachment graphs).
  • On round $1\leq i\leq n$, add a vertex $v_{i}$ into the graph.
  • Then, repeat the following process $\mu$ times:
    • The probability of the edge $(v_{i},v_{j})$ being added is $\frac{d_{v_{j}}}{\sum_{1}^{i}d_{v_{j}}+1}$ if $j\neq i$, and $\frac{d_{v_{j}}+1}{\sum_{1}^{i}d_{v_{j}}+1}$ if $j=i$.
@@ 3 Spanners for Erdős-Rényi Graphs
3.1 Constant Stretch Spanners
See 1.6 We first note that we require a bound on the diameter of random graphs:
@@ 3.1 Constant Stretch Spanners
We next prove a complimentary result, which obtains constant query complexity in a weaker probe model, at the cost of slightly superlinear edge counts. We note that we still obtain a superior sparsity-stretch tradeoff compared to Observation 1.4, as no choice of $p^{\prime}$ can achieve nearly-linear sparsity with constant stretch. See 1.7
Proof.
Our algorithm is the same as Theorem 1.6, except that we modify the connectivity rule to purely depend on the randomness, not the relative index of the edge. First, rather than selecting the center based on vertex IDs, each vertex independently elects itself into the center with probability $2/T$. With high probability, the number of center vertices remains bounded by $[T,3T]$, and hence the number of center internal edges remains $o(n)$, and the stretch of the center subgraph does not increase. Next, to determine if we keep edge $(u,v)\in G$, the connectivity rule is as follows:
  • If $u$ and $v$ elect themselves centers, keep.
  • If $u$ and $v$ do not elect themseselves centers, do not keep.
  • Otherwise (so exactly one endpoint is a center), keep independently with probability $\rho$ for $\rho$ to be chosen later.
It is clear such a connectivity rule can be implemented with $1$ query, to verify $(u,v)$ is actually in the base graph.
Stretch.
Observe that, conditioned on the center graph being of size $[T,3T]$, we can apply the exact argument as before and conclude that the diameter of the center graph is unchanged. Thus, it suffices to show that all non-center vertices remain connected. For every $v$, by essentially the same argument as Theorem 1.6 we have that with high probability
$|\Gamma(v)\cap C|\in[n^{\delta/2-\delta^{2}/8}/2,2n^{\delta/2-\delta^{2}/8}]$
Finally, let $\rho=10\log(n)/n^{\delta/2-\delta^{2}/8}$. We have that with high probability at least one, and fewer than $100\log(n)$ edges are retained from $v$ to the center subgraph, and hence we obtain the desired edge bound. ∎
3.2 Ultra-Sparse 4-Spanners for Dense Random Graphs
For our final result, we first again designate vertices as centers, and choose parameters such that all vertices either connect to a single center, or help two centers connect each other. See 1.8
Proof.
We again apply Observation 1.4, and first subsample $G$ to $G^{\prime}$ with edge probability $p^{\prime}=8\log^{2}(n)/\sqrt{n}$ (which is valid as $\delta>1/2$). For the remainder of the proof, we assume without loss of generality that we have probe access to a graph $G^{\prime}\leftarrow G(n,p^{\prime})$. [^2] We note that in the final algorithm, each time we enumerate the neighbors of $v$, we will pay $np$ queries (rather than $np^{\prime}$).
Global Implementation.
After this, we let the first $\sqrt{n}/\log(n)$ vertices be centers. Denoting this center set as $C$, we have the following connectivity rule. For every pair of centers $v,v^{\prime}$, let $u$ be the smallest vertex such that $(v,u),(v^{\prime},u)\in G^{\prime}$. We retain both edges in $H$. Finally, for every $u\notin C$ that is not connected to a center by the prior rule, let $v$ be the smallest center vertex such that $(u,v)\in G^{\prime}$. We retain this edge in $H$.
To prove the diameter bound, we show that every non-center vertex neighbors a center, and every pair of center vertices share a neighbor:
Claim 3.2.
With high probability, every $v\notin C$ has $|\Gamma(v)\cap C|\geq 1$.
Proof.
The probability that we have no edges from $v$ to $C$ is exactly $1-(1-p^{\prime})^{|C|}\geq 1-e^{p^{\prime}\sqrt{n}/\log n}\geq 1-n^{-4}.$
Claim 3.3.
With high probability, every $v,v^{\prime}\in C$ have $|\Gamma(v)\cap\Gamma(v^{\prime})|\geq 1$.
Proof.
For arbitrary $u\notin C$, the probability that are are edges $(v,u),(v^{\prime},u)$ is $p^{{}^{\prime}2}$. Thus, the probability that there are no vertices $u$ with this property is exactly $1-(1-p^{\prime 2})^{n-|C|}\geq 1-e^{-32\log^{4}(n)}.$
Diameter.
Is it easy to show the diameter is $4$. Given an arbitrary vertex pair $a,b$, both vertices are connected to (possible distinct) centers $v,v^{\prime}$ (or they are themselves centers), and every pair of centers is connected by a path of length $2$.
Sparsity.
We have that the number of included edges is bounded by $2\cdot T^{2}+n$, as there are at most $T^{2}$ length-2 paths (that connect each pair of centers) included, and at most $n$ edges from non-center vertices connecting them to a center vertex.
Local Implementation.
Given an edge query $(a,b)$, we can decide the edge status locally as follows. First, if $(a,b)\notin G^{\prime}$, do not retain it. Otherwise, if $a,b\in C$ or $a,b\notin C$, do not retain the edge. Finally, for the case in which $a\in C$ and $b\notin C$, we enumerate over $u\in\Gamma_{G^{\prime}}(b)$ and, if $u$ is a center vertex, check if $b$ is the smallest vertex in $\Gamma_{G^{\prime}}(u)\cap\Gamma_{G^{\prime}}(a)$. If so, the path $(a,b)(b,u)$ was part of the global connectivity rule, so we retain $(a,b)\in H$. If we do not include any edges for this reason, and $a$ is the smallest center in $\Gamma(b)$, the edge $(a,b)$ was part of the global connectivity rule, so we retain $(a,b)\in H$. Otherwise, we do not retain the edge.
The probes required to implement this rule are bounded as follows. First, we must query the neighborhood of $b$ using $O(n^{\delta})$ probes. Folowing this, we perform the sparsification step from Observation 1.4 to prune down to edges , and then operate on this sparsified graph, which has degree $\widetilde{O}(\sqrt{n})$. To bound the number of center neighbors that we further explore, note that
$|\Gamma_{G^{\prime}}(b)\cap C|=O(p^{\prime}T)=O(\log^{3}n)$
with high probability. For each center neighbor $u$, we must enumerate $\Gamma_{G^{\prime}}(u)$ which takes $O(n^{\delta})$ probes per vertex, so we end up with a final bound of $\widetilde{O}(n^{\delta})$ as claimed. ∎
4 Sparse Connected Subgraphs For Sparse Erdős-Rényi

### v2

@@ 1.2 Our Results: Joint Sampling of Erdős-Rényi Graphs and Maximal Independent Sets
A Local Access Generator (LAG) of a random object $G$ sampled from a distribution $\mathcal{G}$, is an oracle that provides access to $G$ by answering various types of supported queries, given a sequence of random bits $\vec{r}$. We say the LAG is memoryless if it does not store its answers to prior queries. We require that (fixing a random tape) the responses of the local-access generator to all queries must be consistent with a single object $G$. Moreover, the distribution $\mathcal{G}^{\prime}$ sampled by the LAG must be within $n^{-c}$ from $\mathcal{G}$ in TV distance, for any desired constant $c$.
As in the case of LCAs, we desire Local Access Generators to be as efficient as possible per query. We also strongly desire the LAG to be memoryless (a requirement in the setting of LCAs, but not always achieved for LAGs), and our result achieves this goal.
@@
We show that such an approach is also fruitful in the setting of LCAs. We again focus on the dense Erdős-Rényi model, and this time on the extensively studied problem [Gha16, LRY17, GU19, Gha22] of Maximal Independent Set (MIS). The frontier result of Ghaffari [Gha22] provides an LCA for MIS with per-query runtime $\operatorname{poly}(\Delta\log n)$, and a local sampling implementation of dense Erdős-Rényi graphs is straightforward. However, composing these algorithms does not give a sublinear runtime. Our result achieves runtime $\operatorname{polylog}(n)$ for $p\geq 1/\operatorname{polylog}(n)$ per query, both for queries to the random graph and to its MIS:
@@
2 Proof Overviews
2.1 Spanners for Erdős-Rényi
@@
See 1.7 Our connectivity rule is as follows. We designate a sublinear-size set of vertices in $G$ as centers, which we denote $\mathcal{C}$. We then retain in $H$ all edges between centers. Finally, every non-center vertex adds the first edge from itself to $\mathcal{C}$ in $H$.
By choosing the size of $\mathcal{C}$ appropriately, we ensure that the following three conditions hold with high probability: there are $o(n)$ intra-center edges (enforced by choosing the size so that $|\mathcal{C}|^{2}p=o(n)$), every non-center vertex has an edge to the center with high probability (enforced by choosing the size so that $|\mathcal{C}|p=\Omega(\log n)$), and the center has constant diameter (which follows as the center is itself distributed as $G(|\mathcal{C}|,p)$).
To implement this connectivity rule as an LCA, we break into the sparse case (where a non-center vertex simply queries its entire adjacency list and chooses the least-ranked edge to keep) or the dense case (where a non-center vertex queries the adjacency matrix until it finds its first edge). If we additionally assume that the adjacency list of each vertex is sorted in ascending order, we can perform this check in constant time.
2.2 Sparse Connected Subgraphs for Erdős-Rényi
Recall we are given probe access to $G\leftarrow G(n,p)$ and wish to provide local access to a sparse connected subgraph $H\subseteq G$ with very few edges. Here we focus on the case where the input graph is itself somewhat sparse. Without essential loss of generality, we assume the edge probability is exactly $p^{*}=7\log n/n$ (as otherwise we can use the idea of Observation 1.5 to subsample as a first step).
@@ 2.2 Sparse Connected Subgraphs for Erdős-Rényi
Connectivity. We define a set of events $\mathcal{E}$ that partition the space of possible graphs, and denote a subset of events $\mathcal{E}_{G}\subset\mathcal{E}$ as good. We first show that a random graph lies in a good event with high probability. Next, we show that for every good event $E\in\mathcal{E}_{G}$, sampling a random $G$ that satisfies $E$ results in a graph that the algorithm succeeds on (in fact, we prove this with high probability over $G$).
Each such event specifies the presence or absence of a subset of edges in the graph. At a high level, these specifications capture the view of the algorithm up to the point that the leader vertices select their administrators. We define good events as those in which all administrators have many bits of entropy remaining in their neighborhoods, which allows us to argue they maintain connectivity with high probability.
@@
Local Implementation. One can see that the algorithm constitutes a $4$-round distributed algorithm, and hence can be implemented in per-query work $O(\Delta^{4})$ via the reduction of Parnas and Ron [PR07]. However, we note that we can first subsample the graph $G$ to have edge probability $p^{*}=7\log n$ (which we do in a global fashion using the random tape of the LCA). By Observation 1.5, this produces whp a connected subgraph of $G$ that is itself distributed $G(n,p^{*})$. Subsequently, in each distributed round we explore only the neighbors of $v$ that are retained in the subsampled graph, resulting in total work $O(\Delta\log^{4}n)$.
2.3 Spanners for Preferential and Uniform Attachment
We highlight the main ideas behind the proofs of Theorem 1.9 and Theorem 1.10. We are interested in the case where the degree parameter $\mu$ is sufficiently high (and thus the total number of edges $n\cdot\mu$ is high and constructing spanners is compelling). When $\mu$ is large, both preferential and uniform attachment graphs are well-structured, in the sense that higher degree vertices typically have an earlier arrival time. At a high level, we can leverage this structure by having each vertex keep an edge to its highest-degree neighbor, thus generally having the spanner keep paths from vertices to the earliest-added vertices in the process.
For preferential attachment graphs (Theorem 1.9), we use the following algorithm to build low-stretch spanning trees. Because preferential attachment graphs are multigraphs, the algorithm’s input specifies the two nodes adjacent to the edge as well as the lexicographic indexing of the edge. (See Remark 3.1.)
Algorithm: On query $(u,v,i)$:
Check if $v$ is the highest-degree neighbor of $u$ or vice versa. If so, keep the edge if it is the lexicographically first edge between $u$ and $v$ (i.e. $i=1$). Otherwise, discard the edge.
@@ 2.3 Spanners for Preferential and Uniform Attachment
The proof is a direct consequence of a structural result about preferential attachment graphs with $\mu(n)\geq c_{\mu}\log n$: With high probability, every vertex $v$ that is not the highest-degree vertex is either a neighbor of the highest-degree vertex, or $v$ has a neighbor $u$ such that $d_{u}>2\cdot d_{v}$.
@@
To show this structural result, we prove that there is a global constant $c$ such that with high probability, every $v_{i}$ with arrival time $i\geq c$ has a neighbor with degree at least $2d_{v_{i}}$, and every $v_{i},v_{j}$ for arrival times $i,j\leq c$ are connected. The second item is a simple consequence of our choice of $\mu$, and the first follows from a tail bound for the degrees of preferential attachment graphs established by [DKR18]. Given this result, connectivity is direct, and a simple potential argument (that the degree cannot increase by a factor of $2$ more than $\log(m)$ times, where $m$ is the total degree of the graph) establishes a stretch bound of $O(\log(d_{\max}))=O(\log n)$. Furthermore, we show that the worst-case runtime is $O(\mu\sqrt{n})$ and the average-case runtime over all possible queries is $\mu\cdot\operatorname{polylog}(n)$. To see the size bound, note that every vertex that is not of globally highest degree adds exactly one edge to $H$. For a formal proof, see Section 6.
Moving to the setting of uniform attachment graphs (Theorem 1.10), we utilize a similar algorithm that keeps an edge $(u,v)$ if $u$ is the highest-degree neighbor of $u$ or vice-versa. We additionally keep an edge between vertices that have a degree above some threshold (to ensure that edges are kept between the earliest-added vertices, whose degrees will all be similar). The similarity between the algorithms for preferential and uniform attachment demonstrates the robustness of our approach to different settings of randomly growing graph processes. We present the algorithm in Section 7.
We prove that every vertex $v_{i}$ is connected to at least one vertex added sufficiently earlier in the process, and such a vertex can be identified because it will have a higher degree than neighbors added later. The LCA keeps an edge to such a vertex. More formally, we define subsets of vertices $C_{1}\supset C_{2}\supset\dots\supset C_{M}$ for some $M$ that is $O(\log n)$; each $C_{m}$ is defined as the first $|C_{m}|$ vertices added to the process. We show that, with high probability, all vertices in $[n]\setminus C_{1}$ are connected to a vertex in $C_{1}$, and similarly all vertices whose time of arrival $i$ satisfies $|C_{m+1}|<i<e^{2}\cdot|C_{m}|$ are connected to a vertex in $C_{m+1}$, for all $m\in\{1,2,\dots,M\}$. The $e^{2}$ arises because degrees are not precise indicators of times of arrival (see Section 7.3).
Finally, Theorem 1.10 is proven by arguing that, when each vertex keeps an edge to its highest-degree neighbor, the LCA keeps paths of length $O(\log n)$ from each vertex to the vertices added in the first $t$ steps, for a small constant $t$. All adjacencies between the vertices added in the first $t$ steps are preserved, resulting in a low-stretch sparse spanner. For the formal proofs, see Section 7.
2.4 Joint Sampling of Erdős-Rényi Graphs and Maximal Independent Sets
@@
We describe the sampling algorithm in a global fashion, and then describe how to implement it via an LCA. We gradually grow the MIS $M$, by instantiating $M=(1)$ where $1$ is the first vertex, and sequentially determining the smallest vertex that is not connected to $M$, and add it to $M$, continuing until we exhaust the vertex set. For a fixed $M=(v_{1},\ldots,v_{t})$, each subsequent vertex $u>v_{t}$ is not connected to any element of $M$ in $G(n,p)$ with probability $(1-p)^{|M|}$. For a fixed $u$, we can determine if $(v_{i},u)$ is in $G$ for every $i$ by simply sampling the edge, and thus determine if $u$ should be added to $M$. However, this procedure would take linear time to determine the MIS. Instead, we use local sampling of the Geometric distribution to find the next element of $M$. Once we sample $r\sim\mathrm{Geom}(1-p)$, we let the next element of $M$ be $u=v_{t}+r$. For all $u^{\prime}\in(v,u)$, this virtually conditions on the event that at least one of $(u,v_{i})$ is present in the graph for $i\leq t$.
@@ 3 Preliminaries
Remark 3.1 (Remark on multigraphs).
Certain random models we consider will produce multigraphs with high probability. Without loss of generality, we assume that if $\texttt{Nbr}(v,i)=\texttt{Nbr}(v,j)$ for $i<j$ (i.e. there is a multi-edge), the second query returns $\perp$ (alternatively, our algorithm only retains the lexicographically first edge).
@@
We formally define the Erdős-Rényi graph model. Note that in all proofs, $\Gamma(v)=\Gamma_{G}(v)$ refers to the neighbors of $v$ in the original graph $G$ that the LCA has probe access to. We will always denote the original graph as $G$, and subgraphs or spanners as $H$.
@@ 4 Spanners for Erdős-Rényi Graphs
See 1.7 We first note that we require a bound on the diameter of random graphs:
@@
5 Sparse Connected Subgraphs for Erdős-Rényi